# An Algorithmic Approach to Finding Degree-Doubling Nodes in Oriented Graphs

### v10

@@ 1 Introduction
This paper begins by defining the key terminology and tools required to understand the conjecture. We then present our approach, illustrated with lemmas and examples that demonstrate its utility. We then introduce a novel data structure designed to address key challenges, and then we proceed to explore how this data structure can be combined with the decreasing neighborhood sequence property. We examine the role of back arcs and their role in the conjecture. Key subgraph structures, such as transitive triangles and Seymour diamonds, are shown not to impede progress. Our main theorem follows, alongside a detailed analysis of the associated algorithm’s complexity. The paper concludes with potential applications of our approach and insights into future directions for research.
@@ 2 Graph Theory Terminology
This section is for those who are unfamiliar with graph theory. We tried to provide enough background here for readers to follow this paper and the terminology, conjectures, lemmas, theorems, proofs, and algorithms herein.
An oriented graph is a directed graph with no self-loops and no pair of vertices connected by edges in both directions—that is, at most one directed edge exists between any two vertices. The square of a graph $G$, denoted $G^{2}$, is a graph on the same vertex set where an arc exists from $u$ to $v$ if there is a path of length at most two from $u$ to $v$ in $G$. Essentially, $G^{2}$ captures two-step connections from $u$ to $v$.
@@
$N^{+}(v)$ The out-neighbors of a $v$
$N^{++}(v)$ the second out-neighbors of $v$
$d^{+}(v)$ The out-degrees of $v$ in $G$
$d^{+}_{G^{2}}(v)$ The out-degrees of $v$ in $G^{2}$
Figure 2: A five-vertex oriented graph representing a cycle. We can see that every node
@@
Figure 3: A five-vertex oriented graph showing original edges with the full edges and the square cycles with the dashed edges. What we can see here is that every node satisfies both versions of the conjecture.
We can illustrate the equivalence between Conjectures 1.1 and 2.1 by showing that the same nodes satisfy the degree-doubling condition in both cases. To do so we will consider a graph $G$ that consists of a cycle on $n$ nodes and $n$ arcs, represented as $v_{0}\to v_{2}\to v_{3}\to\dots\to v_{n-1}\to v_{0}$
First, we will consider Conjecture 1.1. Let $v_{i}$ be a node on this cycle. Then we know its first out-neighbors $N^{+}(v_{i})=\{v_{i+1}\}$ (Indices are modulo $n$). Its set of second out-neighbors is $N^{++}(v_{i})=\{v_{i+2}\}$. Thus, we have that $|N^{++}(v_{i})|=1$ and $|N^{+}(v_{i})|=1$. This means that there are at least as many second out-neighbors as first out-neighbors, and every node $v_{i}$ in the cycle satisfies the requirements of the conjecture ($|N^{++}(v_{i})|\geq|N^{+}(v_{i})|$).
Next, we consider the same cycle in Conjecture 2.1. For every node $v_{j}$ in the directed cycle, the out-degree in $G$ is $d^{+}_{G}(v_{i})=1$. In the squared graph $G^{2}$, there are arcs $(v_{i},v_{i+1})$ (modulo n) as well as $(v_{i},v_{i+2})$, so the out-degree is $d^{+}_{G^{2}}(v_{i})=2$. Thus we have $d^{+}_{G^{2}}(v_{i})\geq 2\cdot d^{+}_{G}(v_{i})$ in this case.
@@
3 Exploring the Contrapositive
Our approach to the Seymour Second Neighborhood Conjecture was driven by a desire to explore the contrapositive. We did not directly search for a counterexample. Instead, we aimed to understand the implications of assuming the conjecture is false. As data scientists, we naturally gravitated towards an exploratory methodology, programming tools to observe patterns, features, and potential contradictions arising from this assumption. This approach allows us to define concrete conditions that, if violated, directly lead to the identification of a degree-doubling node through our algorithmic process.
One fundamental question we must answer is why we chose proof by contradiction when so few others who have investigated this problem did so. First, from a practical point of view, it is necessary to use all tools at our disposal. Proof by contradiction is one of those tools and is logically equivalent to alternative proof methods. As computer programmers, though, what proof by contradiction does is often re-frames the problem from a theoretical point of view into an actionable one. That is, the proof by contradiction gives us that extra assumption that often allows for datasets and concrete conditions that can guide the development of an algorithm. Finally, we recognize that fresh perspectives on a long-standing problem are difficult to come by. Investigating the contrapositive offers the insight to explore less-traveled paths and bring about new proofs by simply taking ”The Road Less Traveled”.
For the Second Neighborhood Conjecture (SSNC), an additional insight came both from intuition—shaped by our work on the Directed Second Neighborhood Property (DSNP)—and from a new structural feature: exterior neighbors. While interior neighbors (which align with transitive triangle structures) have appeared in the literature in various forms, exterior neighbors have largely gone unmentioned. Yet these two concepts are fundamentally linked. They are, in essence, two sides of the same coin—capturing local and extended connectivity within a graph. Since the SSNC concerns both first and second out-neighbors, any attempt to resolve or even approach the conjecture rigorously must treat both interior and exterior neighbors in parallel. One cannot be fully understood or leveraged without the other. What we will see is that this extra data point is allowed to build a traversal through the graph. These exterior arcs connect different partitions, or nodes at a distance from a chosen root node.
Early in the research process, we began to notice that the SSNC places two fundamental yet distinct graph metrics in competition: distance and degree. The conjecture seeks for a relationship between the size of a node’s first out-neighborhood, which is based on direct distance, and its second out-neighborhood, which is based on a distance of two. However, the question the conjecture asks is for a node whose degree doubles. This shows two metrics that are intertwined into one conjecture. Other graph metrics such as centrality or tree width could be considered, but they are not intrinsic to the conjecture. Degree and distance, however, are explicitly central, and any resolution of the SSNC must contend with their relationship.
When we were faced with the question of which metric to prioritize for our initial exploration and partitioning strategy, both metrics gave reasonable options, but for different reasons. For example, every node has an out-degree, and thus there exists a minimum value of these out-degrees. This is the out-degree value that all other nodes in the graph must at least have. This was a reasonable starting point for exploratory analysis. In the SSNC, distance has also received a lot of attention. There is the issue of double-counting, where second out-neighbors are defined in terms of reachability (i.e., they can be members of $N^{+}(x)$ or $N^{++}(x)$), or defining second out-neighbors in terms of disjoint union to make sure that this double-counting issue is addressed, but potentially losing track of paths. We chose to focus our attention on distance with an anchor node as the minimum out-degree node. This allows us to partition the graph about that node and take advantage of some of the other properties of distance, such as total order. This will be defined more formally in Section 5 (Graph Level Order). As such, this will keep track of paths in our partitions, unlike previous approaches.
The next step in our analysis is the consideration of arbitrary oriented graphs. We did not want to make assumptions about node degrees or other graph properties that were too restrictive. For example we did not want to assume, at least initially, that any class of subgraphs were excluded from these oriented graphs. Instead, we wanted to consider the fundamental characteristics common to all oriented graphs. As we mentioned before, we know that there is an out-degree distribution. Consequently, every oriented graph will have a minimum out-degree $\delta$, representing the lowest out-degree among all nodes. By selecting a node $v_{0}$ with this minimum out-degree, we establish a stable starting point for our investigation.
The notion of partitioning the graph into subgraphs is not introducing the NP-hard problem of set partitioning into the SSNC. The set-partition problem asks if a sequence of positive numbers $S$ can be partitioned into two sets $S_{1},S_{2}$ such that the sum of the numbers in those sets are equal. That is a search for a partition. On the contrary, this algorithm begins with an anchor about the minimum out-degree node and declares that the partition be done based on that node.
  • Shared First Out-Neighbors $N^{+}(u)\cap N^{+}(v)$: $u$ is a node that is closer to $v_{0}$ than $v$, and they share common first out-neighbors. These are the shared direct out-neighbors of $u$ and $v$. This is the situation where $v$ assists with $u$ not being a degree-doubling node.
  • $u$’s Second vs. $v$’s First $N^{++}(u)\cap N^{+}(v)$: $u$ is a node that is closer to $v_{0}$ than $v$, and $u$’s second out neighbors are common with $v$ first. These are the out-neighbors of $v$ that are reached in two steps from $u$. This gives rise to $u\rightarrow y_{1}\rightarrow x$ and $v\rightarrow x$. We have two cases. Either $y_{1}\in N^{+}(u)$ or $y_{1}\notin N^{+}(u)$. The first case is $y_{1}\in N^{+}(u)$. Then $N^{+}(u)\cap N^{+}(v)\neq\emptyset$ (because of $y_{1}$). The second case is $y_{1}\notin N^{+}(u)$. Then $N^{++}(u)\cap N^{+}(v)\neq\emptyset$ (again because of $y_{1}$).
  • $u$’s First vs. $v$’s Second $N^{+}(u)\cap N^{++}(v)$: These are the out-neighbors of $u$ that are reached in two steps from $v$. This implies that we have $u\rightarrow x$ and $v\rightarrow y_{1}\rightarrow x$. Remember that $v\in N^{+}(u)$, which means that we have $u\rightarrow v$. So $x$ and $v$ have the same distance from $v_{0}$. Let $I=N^{+}(u)\cap N^{++}(v)$. If there exist a $z\in I$, then $z\in N^{+}(u)$ and $z$ has the same distance from $v_{0}$ as $x$ and $v$. This implies that the arcs of the form $N^{++}(v)$ do not increase the distance of nodes in $I$.
  • Shared Second Out-Neighbors $N^{++}(u)\cap N^{++}(v)$: These are the common second out-neighbors of both $u$ and $v$. This would mean that they are second out-neighbors within the first out-neighborhood of $u$. Remember, though that $v\in N^{+}(u)$, so we have two paths to a node $w$. $u\rightarrow y_{1}\rightarrow w$ and $u\rightarrow v\rightarrow z_{1}\rightarrow w$. This says that $w$ is reachable from both length 2 and length 3 from $u$. Similar to the case $N^{+}(u)\cap N^{++}(v)$, this is dependent on $N^{++}(u)\cap N^{+}(v)$ because without the arc $(v,y_{1})$ in this example, there is no path from $v$ to $w$, and so $N^{++}(u)\cap N^{++}(v)$ would be $\emptyset$.
These four intersection types allow us to track how influence and load-sharing propagate from $u$ through its neighborhood, giving a nuanced lens on why certain nodes fail to double their degree. We see from these four cases that the main two cases that need to be concentrated on are the first two, namely when there are shared first out-neighbors, and shared second with first. The rest of this paper will delve more into these cases.
As we noted earlier, this proof by contradiction begins at the node level. We are constructively building this house of cards with our definitions, lemmas, examples, theorems and ultimately our algorithms. Some of the first questions we will ask are what things need to be in place in a graph in order for a node’s degree not to double. We will begin this process with minimum degree nodes, and branch out from there. At the same time, these lemmas we are constructing are helping the house of cards to grow collectively stronger, to ultimately either produce a degree doubling node or prove that one cannot exist.
We will revisit the concept of transitive triangles, not as hindrances but as mechanisms for load sharing. These triangles allow for the neighbors of a node to support one another. This can help prevent their parent’s degrees from doubling. These shared responsibilities can prevent an immediate predecessor from being a Seymour vertex or not having their degrees double. The question then becomes: how much order is embedded in these graphs, and how frequently do these patterns emerge? Consider the example of a node $u$ with three out-neighbors $v_{1},v_{2},v_{3}$ such that we have a cycle in the subgraph $v_{1}\rightarrow v_{2}\rightarrow v_{3}\rightarrow v_{1}$. This cycle prevents node $u$’s degree from doubling. This is $u$ sharing its load with its first out-neighbors.
We must also discuss complexity when discussing global search algorithms. This raises concerns about how well they will handle larger graphs in comparison to smaller ones. This should not be a problem, as we will see in the following sections. The initial step of our algorithm involves partitioning the graph’s nodes and arcs based on their distance from the minimum out-degree node, which can be done in $O(V+E)$ time using Dijkstra’s Algorithm which returns a vector containing the lengths of the shortest paths to all nodes in the graph. What we will then see is that these partitions decrease in size as we move from the minimum out-degree node. Our algorithm will simply take a path from the minimum out-degree node to the degree-doubling node.
Finally, we turn our attention to the term degree-doubling nodes. By this term, we aim to get to the heart of Conjecture 2.1. That conjecture is looking for a node whose degree at least doubles in the square of the graph. By the equivalence of the two conjectures, this is the same as looking for a node whose second neighborhood is at least as large as its first, i.e. a node that satisfies Conjecture 1.1. A similar definition, a Seymour vertex, fulfills the conjectures—it’s out-degree in the squared graph is at least double its original out-degree. We could equivalently search for a Seymour vertex or a degree-doubling node. We note that we can make a degree-doubling node a noun, a property of the node having its degree double, or a verb, as in having its degree doubled. We do not have such freedom with the term Seymour vertex.
Definition 3.1.

### v12

@@ 1 Introduction
This paper begins by presenting the methodology behind our approach. We then define the key terminology and tools required to understand the conjecture. We follow that with lemmas and examples that demonstrate its utility. We then introduce the novel data structure designed to address key challenges, and then we proceed to explore how this data structure can be combined with the Decreasing Neighborhood Sequence Property. We examine the role of back arcs and their role in the conjecture. Our main theorem follows, alongside a detailed analysis of the associated algorithm’s complexity. The paper concludes with potential applications of our approach and insights into future directions for research.
2 Exploring the Contrapositive
Our approach to SSNC was driven by a desire to explore the contrapositive. We were not directly searching for a counterexample. Instead, we were aiming to understand the implications of the conjecture, and what it would mean to assume it false. As data scientists, we naturally gravitated towards an exploratory methodology, programming tools to observe patterns, features, and potential contradictions arising from this assumption. This approach allows us to define concrete conditions that, if violated, directly lead to the identification of a degree-doubling node through our algorithmic process.
One fundamental question we must answer is why we chose proof by contradiction when so few others who have investigated this problem did so. First, from a practical point of view, it is necessary to use all tools at our disposal. Proof by contradiction is one of those tools and is logically equivalent to alternative proof methods, like direct proof. As computer programmers, though, what proof by contradiction does is re-frames the problem from a theoretical point of view into an actionable one. That is, the proof by contradiction gives us an extra assumption that allows us to construct datasets and concrete conditions that can guide the development of an algorithm or insight. Finally, we recognize that fresh perspectives on a long-standing problem are difficult to come by. Investigating the contrapositive offered the insight to explore less-traveled paths and bring about new proofs by simply taking ”The Road Less Traveled”. This offers a higher probability of a breakthrough by simply avoiding the obstacles that previous authors have faced, like complexity issues.
For the Seymour Second Neighborhood Conjecture (SSNC), an additional insight came both from intuition—shaped by our work on the Decreasing Neighborhood Sequence Property (DNSP), which is a direct negation of the conjecture but at the node level—and from a new structural feature: exterior neighbors. While interior neighbors (which are a particular type of transitive triangle) have appeared in the literature in various forms, exterior neighbors have largely gone unnoticed. Yet these two concepts are fundamentally linked. They are, in essence, two sides of the same coin—capturing local and extended connectivity within a graph. The SSNC concerns both first and second out-neighbors, so any attempt to resolve the conjecture rigorously must treat both interior and exterior neighbors in parallel. We cannot fully understand one without the other. Interior neighbors are neighbors shared with a parent from the previous neighborhood within and another member of the same neighborhood. Exterior neighbors, on the other hand, are neighbors between a parent and child in consecutive neighborhoods that are not shared by the parent. This will be defined formally in Section 5 (Graph Level Order). What we will see is that this extra data point allows us to build a traversal algorithm through the graph. These exterior arcs connect to the second neighbors of the parent node, while the interior arcs connect to the first neighbors of the parent node.
2.1 A Programmer’s Insight
Early in the research process, we began to notice that the SSNC places two fundamental yet distinct graph metrics in competition: distance and degree. The conjecture seeks for a relationship between the size of a node’s first out-neighborhood, which is based on direct distance, and its second out-neighborhood, which is based on a distance of two. However, the question the conjecture asks is for a node whose degree doubles. This shows two metrics that are intertwined into one conjecture. Other graph metrics such as centrality or tree width could be considered, but they are not intrinsic to the conjecture. Degree and distance are explicitly stated, and any resolution of the SSNC must contend with their relationship.
When we were faced with the question of which metric to prioritize for our initial exploration and partitioning strategy, both metrics gave reasonable options, but for different reasons. For example, every node has an out-degree, and thus there exists a minimum value of these out-degrees. This is the out-degree value that all other nodes in the graph must at least have. This was a reasonable starting point for exploratory analysis. On the contrary, the distance metric has properties like a total order that could prove useful.
Once we had chosen to start with a minimum out-degree node, our approach remained largely procedural, guided by small, illustrative examples. Before we partitioned the graph, even small graphs offered little insight into the graph’s structure. The main theme of proof by contradiction is that we did not want any node’s degree to double. This is what guided our early exploratory process through the graph. Partitioning the graph by distance from the out-degree of the root node (by $\delta$) was not a goal. Instead, we observed that, much like we will see in Examples 4.1 and 4.2, the nodes were naturally grouping themselves into rooted neighborhoods based on their distance from the anchor node. This partitioning was not imposed, it emerged. This became a foundational insight: by organizing nodes in this manner, into these distance-based layers, we gained a clearer view of their roles and dependencies. The partitioning helped us to understand interior flow, exterior flow and identify patterns. This eventually helped us to formalize a framework in which degree doubling could be analyzed systematically.
The first observation we see from Figure 1 is numerical is that these neighborhoods are decreasing in size. We wondered if that was a coincidence, or a consequence of the graph’s local shape. Or did it hold more broadly. This is a proof by contradiction, so the reason for that in this example is because nodes cannot have their second neighborhoods larger than their first neighborhoods. We called this the Decreasing Neighborhood Sequence Property (DNSP) for this reason. Then we notice something more interesting: the correlation of the interior out-degree (the blue arcs) with the neighborhood distance from the root node. Each node in the first neighborhood has an interior out-degree of one. Then each node in the second neighborhood has an interior out-degree of two. A natural question was arising, are these out-degrees typical, or exceptions? If they were typical, can we leverage them towards finding a degree doubling node?
Another pattern that we noticed was that some nodes had first neighbors within the same neighborhood, sharing a common parent. This formed our concept of interior neighbors. Other nodes had neighbors in the next neighborhood, that were the parent’s second neighbor. These are the nodes we identified as exterior neighbors. This was not just a cosmetic difference. We could partition a node’s degree into interior degree and exterior degree, unlike first neighbors and second neighbors. This helps to uncover the mechanics of load sharing, what is necessary for a node to not be a degree doubling node in a graph. It also was not just present in the first partition, but we were seeing it throughout the graph. Interior neighbors represented the interlocking within the neighborhood, while exterior neighbors represented forward pressure. Recognizing and formalizing this split became essential for understanding how rooted neighborhoods either sustain themselves or collapse under the absence of degree doubling nodes. More importantly were the questions of whether these observations were simple examples, or could they be backed up with mathematical rigor.
Figure 1: This is an early visualization of the SSNC through JavaScript’s Canvas. Nodes are colored based on their distance from the node 00. Interior arcs are blue, while exterior arcs are gray. We see that there are seven nodes in the first rooted neighborhood, matching the degree of node 0. There are six nodes in the second rooted neighborhood, and five nodes in the third rooted neighborhood. Each node in the first rooted neighborhood has degree one. Each node in the second rooted neighborhood has degree two. Each Node in the third rooted neighborhood cannot have degree three because ($5\choose 2$ is 10, which is the complete graph on 5 nodes, and each node having degree three would require 15 nodes.
This is not a simple Breadth First Search (BFS) algorithm. Where BFS ends is where the insight begins. First, we notice that the sizes of these neighborhoods are decreasing. That is not simply because of BFS. If we had selected a different graph, with non-decreasing neighborhoods, what we would have seen is a trivial degree-doubling node. What stopped this was that proof by contradiction assumption and the Decreasing Neighborhood Sequence Property (DNSP). When graphs do not have this property, they admit degree doubling nodes. Such trivial cases have been omitted here for simplicity, but BFS would not ignore them. Since we are in a proof by contradiction, BFS trivially producing a degree doubling node because of the structure of the graph, is not what we want. Instead we need the DNSP to constrain the neighborhood growth and require nodes to have the size of their second neighbors to be strictly less than that of their first.
Secondly, These neighborhoods act as node level identifiers, which can first help with distinguishing between different arc types. This is where the notion of interior neighbors, exterior neighbors and back arcs comes from. Further, we can use these same node level identifiers to classify triangles into six distinct types based on the endpoints of the nodes. Ultimately, we will use this classification of arcs into interior, exterior or back as one of the main resources to utilize to prove the truth of the SSNC.
A contrary but clarifying idea also began to take shape regarding the distance metric. Some authors define second out-neighbors purely in terms of reachability - a node belongs to $N^{++}(x)$ if it is reachable via a directed path of length two. There is the issue of double-counting, where a node can be a member of $N^{+}(x)$ or $N^{++}(x)$. Other authors attempt to resolve this double counting by defining second out-neighbors as the disjoint union, but doing so potentially loses track of paths that we would like to reconstruct.
If this representation could lead to a rigorous mathematical foundation, then it has the potential to both keep track of paths, and not have the problems of defining second neighbors in terms of reachability instead of shortest path. The mathematical formulation for these ’rooted’ neighborhoods is simply the shortest path distances from the chosen minimum out-degree node $v_{0}$ to each node in that neighborhood, which is the root node. This allows us to partition the graph about that node and take advantage of some of the other properties of distance, such as total order. This will be defined formally in Section 5 (Graph Level Order). As such, this will keep track of paths in our partitions, unlike previous approaches.
These initial steps were driven by intuition. These small, preliminary examples all identified degree-doubling nodes in highly predictable, structured circumstances. This does not suffice for mathematical rigor. These graphs can be deeply deceptive, as what appears obviously true or general in these limited settings may not hold under greater scrutiny. More difficult questions quickly arose like how do these graphs perform under the presence of more complex back arcs. Such arcs disrupt the interior and forward flow of arcs between rooted neighborhoods. Could we trust this visual intuition? Would these examples scale to arbitrary graphs? More importantly, how can we build a mathematical framework that generalizes this phenomena we observed? The remainder of this work is devoted to answering those questions with rigor.
Indeed, in 2020, we were confident enough in an early version of this framework to share a preliminary write-up with Dr. Paul Seymour at Princeton University. He graciously responded and pointed out a key structural issue: we had missed a case in the presence of back arcs. These are arcs directed from nodes in later neighborhoods back to earlier ones. The presence of these arcs would cause the neighborhoods to overlap, invalidating our initial attempts to treat the sets as disjoint for our inductive for additive arguments.
While our small constructed examples showed that such back arcs often quickly led to a degree doubling node, Dr. Seymour’s example was more abstract but grounded in the language of our definitions. His insight revealed that the real obstacle was not the back arcs themselves, or even the transitive triangles in which they presented themselves. We needed to prove that what we were seeing in larger, more concrete examples, would hold up in this more abstract one. This would take a more structured environment than the one we were working with. This feedback led directly to the development of the Graph Level Order data structure, which was a way to retain order, enforce layered progression, and build the foundation necessary for an algorithmic approach to the SSNC.
2.2 Set Theoretic Aspects
The next step in our analysis is the consideration of arbitrary oriented graphs. We did not want to make assumptions about node degrees or other graph properties that were too restrictive. For example we did not want to assume, at least initially, that any class of subgraphs were excluded from these oriented graphs. Instead, we wanted to consider the fundamental characteristics common to all oriented graphs. As we mentioned before, we know that there is an out-degree distribution. Consequently, every oriented graph will have a minimum out-degree, $\delta$, representing the lowest out-degree among all nodes. By selecting a node $v_{0}$ with this minimum out-degree, we establish a stable starting point for our investigation.
The notion of partitioning the graph into subgraphs is not introducing the NP-hard problem of set partitioning into the SSNC. The set-partition problem [13] asks if a sequence of positive numbers $S$ can be partitioned into two sets $S_{1},S_{2}$ such that the sum of the numbers in those sets are equal. That is a search for a partition. On the contrary, this algorithm begins with an anchor about the minimum out-degree node and declares that the partition be done based on that node.
  • Shared First Out-Neighbors $N^{+}(u)\cap N^{+}(v)$:
    $u$, the parent node, is a node that is closer to $v_{0}$ than $v$, the child node, and they share common first out-neighbors. These are the shared direct out-neighbors of $u$ and $v$. This is the situation where $v$ assists with $u$ not being a degree-doubling node. This is the situation we refer to as interior neighbors.
  • $u$’s Second Neighbors vs. $v$’s First Neighbors Let $x\in N^{++}(u)\cap N^{+}(v)$:
    $u$, the parent node, is a node that is closer to $v_{0}$ than $v$, the child node, and $u$’s second out neighbors are common with $v$ first. These are the out-neighbors of $v$ that are reached in two steps from $u$. This gives rise to $u\to y_{1}\to x$ and $v\to x$. This is the situation where $x$ is reachable through $y_{1}$ or $v$. This illustrates what [2] defined as a Seymour diamond. Since these paths are traceable we do not lose count of the neighbors of $u$ in doing this. This is the situation we refer to as exterior neighbors.
  • $u$’s First Neighbors vs. $v$’s Second Neighbors $x\in N^{+}(u)\cap N^{++}(v)$:
    $u$, the parent node, is a node that is closer to $v_{0}$ than $v$, the child node and these are the out-neighbors of $u$ that are reached in two steps from $v$. This implies that we have $u\to x$ and $v\to y_{1}\to x$. Remember that $v\in N^{+}(u)$, which means that we have $u\to v$. So $x$ and $v$ have the same distance from the minimum out-degree node. This implies that the nodes $x,y_{1},z$ are all the same distance from the minimum out-degree node $v_{0}$. There are two cases to consider If $y_{1}\in N^{+}(u)$ or $y_{1}\notin N^{+}(u)$. If $y_{1}\in N^{+}(u)$ then $y_{1}\in N^{+}(u)\cap N^{+}(v)$. Similarly, if ($y_{1}\notin N^{+}(u)$, then $y_{1}\in N^{++}(u)\cap N^{+}(v)$.
  • $u$’s Second Neighbors vs $v$’s Second Neighbors $x\in N^{++}(u)\cap N^{++}(v)$:
    $u$, the parent node, is a node that is closer to $v_{0}$ than $v$, the child node and these are the common second out-neighbors of both $u$ and $v$. This would mean that they are second out-neighbors within the first out-neighborhood of $u$. Consider the node $z_{1}$. There are two possible cases. Either $z_{1}\in N^{+}(u)$ or $z_{1}\notin N^{+}(u)$. If $z_{1}\in N^{+}(u)$, then $z_{1}\in N^{+}(u)\cap N^{+}(v)$. Similarly, if $z_{1}\notin N^{+}(u)$, then $z_{1}\in N^{++}(u)\cap N^{+}(v)$.
These four intersection types allow us to track how influence and load-sharing propagate from $u$ through its neighborhood, giving a nuanced lens on why certain nodes fail to double their degree. We see from these four cases that the main two cases that need to be concentrated on are the first two, namely when there are first out-neighbors (interior neighbors), and when a parent shares second neighbors with a child’s first neighbors (exterior neighbors). The rest of this paper will delve more into these cases.
As we noted earlier, this proof by contradiction begins at the node level. We are constructively building this house of cards with our definitions, lemmas, examples, theorems and ultimately our algorithms. Some of the first questions we will ask concern what conditions must be in place in a graph in order for a node’s degree not to double. We will begin this process with minimum degree nodes, and branch out from there. At the same time, these lemmas we are constructing are helping the house of cards to grow collectively stronger, to ultimately either produce a degree doubling node or prove that one cannot exist.
We will revisit the concept of transitive triangles, not as hindrances but as mechanisms for load sharing. These triangles allow for the neighbors of a node to support one another. This can help prevent their parent’s degrees from doubling. These shared responsibilities can prevent an immediate predecessor from being a Seymour vertex or not having their degrees double. The question then becomes: how much order is embedded in these graphs, and how frequently do these patterns emerge? Consider the example of a node $u$ with three out-neighbors $v_{1},v_{2},v_{3}$ such that we have a cycle in the subgraph $v_{1}\to v_{2}\to v_{3}\to v_{1}$. This cycle prevents node $u$’s degree from doubling. This is $u$ sharing its load with its first out-neighbors.
We will also see that node’s degree can be partitioned into three distinct components: interior neighbors, exterior neighbors, and a third called back arcs. This represents a clear distinction from standard formulations of the SSNC, where most authors consider only first and second neighbors-without a structural partitioning of the degree. We already illustrated this in Figure 1. As we will see throughout this paper, the partitioning of a node’s degree, in conjunction with the DNSP, plays a crucial role in determining how neighborhood sizes evolve as we move farther from the root node.
Complexity must be addressed when discussing global search algorithms. This raises concerns about how well they will handle larger graphs in comparison to smaller ones. This should not be a problem, as we will see in the following sections. The initial step of our algorithm involves partitioning the graph’s nodes and arcs based on their distance from the minimum out-degree node, which can be done in $O(|E|+|V|)$ time using the Breadth First Search Algorithm (BFS) which returns the levels, corresponding to the rooted neighborhoods. What we will then see is that these partitions decrease in size as we move from the minimum out-degree node. Our algorithm will simply take a path from the minimum out-degree node to the degree-doubling node.
Finally, we turn our attention to the term degree-doubling nodes. By this term, we aim to get to the heart of Conjecture 3.1, which will be introduced in Section 3 (Graph Theory Terminology). That conjecture is looking for a node whose degree at least doubles in the square of the graph. By the equivalence of the two conjectures, this is the same as looking for a node whose second neighborhood is at least as large as its first, i.e. a node that satisfies Conjecture 1.1. A similar definition, a Seymour vertex, fulfills the conjectures—it’s out-degree in the squared graph is at least double its original out-degree. We could equivalently search for a Seymour vertex or a degree-doubling node. We note that we can make a degree-doubling node a noun, a property of the node having its degree double, or a verb, as in having its degree doubled. We do not have such freedom with the term Seymour vertex.
Definition 2.1.
In an oriented graph $G$, a node $v$ is a degree-doubling node (or Seymour vertex) if $|N^{++}(v)|\geq|N^{+}(v)|$, where $N^{+}(v)$ and $N^{++}(v)$ denote the first and second out-neighborhoods of $v$ in $G$, respectively.
@@ 3 Graph Theory Terminology
This section is intended for readers who may not have a traditional background in graph theory. In particular, graph theory is an interdisciplinary field, finding applications in mathematics, computer science, engineering, biology, and social sciences and this proof of the SSNC has reached many of those fields. To make this paper more accessible to a broader audience, we tried to provide a concise overview of some of the key definitions and concepts. We understand that seasoned graph theorists may be familiar with much of the language used throughout, but even the concepts of oriented graphs and square graphs can span beyond a typical graph theory syllabus. Because of the nature of oriented graphs, we write $u\to v$ to indicate a directed arc from $u$ to $v$. This should help with the over-usage of parenthesis in functions and other areas of graph theory. We hope this section serves as a helpful reference point throughout the paper.
An oriented graph is a directed graph with no self-loops and no pair of vertices connected by edges in both directions—that is, at most one directed edge exists between any two vertices. The square of a graph $G$, denoted $G^{2}$, is a graph on the same vertex set where an arc exists from $u$ to $v$ if there is a path of length at most two from $u$ to $v$ in $G$. Essentially, $G^{2}$ captures all two-step connections from $u$ to $v$.
@@
A directed graph $G$ is called oriented if it has no self-loops (i.e., no arcs of the form $u\to u$ where $u$ is a node in $G$) and no symmetric arcs, that is, no arcs of the form $u\to v$ and $v\to u$ where $u$ and $v$ are nodes in $G$.
@@
The second out-neighborhood of a vertex $v\in G$ is defined as
$N^{++}(v)=\{u\in G\mid\exists w\in G\text{ such that }v\to w\to u,\text{ and }u\notin N^{+}(v)\}$
@@
Figure 3: A five-vertex oriented graph showing original edges with the full edges and the square cycles with the dashed edges. Every node satisfies both versions of the conjecture since they all have out-degree of 1.
We can illustrate the equivalence between Conjectures 1.1 and 3.1 by showing that the same nodes satisfy the degree-doubling condition in both cases. To do so we will consider a graph $G$ that consists of a cycle on $n$ nodes and $n$ arcs, represented as $v_{0}\to v_{2}\to v_{3}\to\dots\to v_{n-1}\to v_{0}$. We can see this through Example 3.2 with a graph on five nodes.
First, we will consider Conjecture 1.1. Let $v_{i}$ be a node on this cycle. Then we know its first out-neighbors $N^{+}(v_{i})=\{v_{i+1}\}$ (Indices are modulo $n$). Its set of second out-neighbors is $N^{++}(v_{i})=\{v_{i+2}\}$. Thus, we have that $|N^{++}(v_{i})|=1$ and $|N^{+}(v_{i})|=1$. This means that there are at least as many second out-neighbors as first out-neighbors, and every node $v_{i}$ in the cycle satisfies the requirements of the conjecture $|N^{++}(v_{i})|\geq|N^{+}(v_{i})|$.
Next, we consider the same cycle in Conjecture 3.1. For every node $v_{j}$ in the directed cycle, the out-degree in $G$ is $d^{+}_{G}(v_{i})=1$. In the squared graph $G^{2}$, there are arcs $v_{i}\to v_{i+1}$ (modulo n) as well as $v_{i}\to v_{i+2}$, so the out-degree is $d^{+}_{G^{2}}(v_{i})=2$. Thus we have $d^{+}_{G^{2}}(v_{i})\geq 2\cdot d^{+}_{G}(v_{i})$ in this case.
@@
For readers seeking a more formal or comprehensive introduction to graph theory, we recommend standard references such as Reinhard Diestel’s Graph Theory [10] provides a thorough and widely used introduction to the standard foundations of the field. Graphs, Networks and Algorithms [19] by Dieter Jungnickel approaches the subject from a computer science perspective.